Roja swathi please explain how to attend the motion of train problems. Que problems1-a train 280 mts long is running at a speed of 42km/hr.How much time will it take to pass a man standing on a platform. Que problem 2 a train 350 mtrs long crosses an electric pole in 28 secs. find the speed of the train in km/hr please explain clearly because i am a slow understander.

Roja  swathi please explain
how to attend the motion of train problems.
 
Que problems1-a train 280 mts long is running at a speed of 42km/hr.How much time will it take to pass a man standing on a platform.
 
Que problem 2 a train 350 mtrs long crosses an electric pole in 28 secs. find the speed of the train in km/hr
 
please explain clearly because i am a slow understander.

Grade:8

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
Convert Speed of train intom/s 42km/hr= 42*5/18 m/s =35/3 m/s
Time taken to cross men will length of train/speed of train = 280/35/3 = 24sec.

For getting speed it will length of train/time = 350/28 m/s for speed in km/hr it will be 350*18/28*5 =90km/hr